OpenStudy (anonymous):

How do you find the indefinite integral of [(1+e^x)^2]/e^x?

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

So you have: \[\int\limits \frac{(1+e^x)^2}{e^x}dx\]

OpenStudy (anonymous):

\[\int\limits \frac{(1+e^x)^2}{e^x}dx\]\[\int\limits\frac{1+2e^x+e^{2x}}{e^x}dx=\int\limits e^{-x}dx+\int\limits 2dx +\int\limits e^xdx\]

OpenStudy (anonymous):

Multiply the top out. Then divide the e^x through leaving: \[\int\limits \frac{1+2e^x+e^{2x}}{e^x}dx=\int\limits [e^{-x}+2+e^x]dx\].

OpenStudy (anonymous):

-e^x+2x+e^x=2x+2sinhx

OpenStudy (anonymous):

Then from there its straight forward: I=-e^-x+2x+e^x+c=2x+2sinh(x)+c

OpenStudy (anonymous):

Haha, at least I wasn't the only one noticing the hyperbolic trig :D

OpenStudy (anonymous):

Make sense blue?

OpenStudy (anonymous):

I get that, thanks :D

OpenStudy (anonymous):

:D